# Painlevé solitons of AKNS system and irrational algebraic solitons of NLS equations

### v1

@@ 1 Introduction
The study of integrable nonlinear partial differential equations (PDEs) has been a central theme in mathematical physics for over half a century. Among the most influential discoveries in this field is the inverse scattering transform (IST), a powerful method for solving certain nonlinear evolution equations. A major breakthrough in the systematic development of the IST was the formulation of the Ablowitz-Kaup-Newell-Segur (AKNS) system [1]. This elegant framework provides a unified approach to treat a wide class of integrable equations, including the celebrated nonlinear Schrödinger (NLS), Korteweg-de Vries, and sine-Gordon equations, by associating them with a compatible pair of linear differential equations (a Lax pair). Within this scheme, the search for special solutions, particularly solitons, and the investigation of their interactions have revealed profound connections with other branches of mathematics, such as the theory of special functions and Painlevé transcendents.
@@
A significant step in this direction is the concept of the “elliptic solitons” [2]. Here, the soliton propagates on a background given by an elliptic (or cnoidal) wave, which itself is an exact periodic solution of the NLS equation. The elliptic wave is expressed in terms of Jacobi elliptic functions. The composite solution, the elliptic solitons, can be derived through sophisticated methods such as the Darboux transformation or Bäcklund transformation, applied to the finite-gap (periodic or quasi-periodic) solutions of the integrable system. The existence of such solutions is deeply tied to the underlying symmetries of the equation. Specifically, it has been shown that the combination of the space-time translation invariance and the so-called “square eigenfunction symmetry” of the AKNS system is instrumental in generating these elliptic soliton solutions.
@@
where $\alpha$ and $\beta$ are complex parameters, emerges naturally in similarity reductions of the NLS equations, the derivative NLS equations, the Boussinesq equation and other integrable systems. Its solutions, the Painlevé IV transcendent, provide a rich class of potential backgrounds that are asymptotically rational, algebraic, or expressible in terms of parabolic cylinder functions and error functions.
By analogy with elliptic solitons, one can define “Painlevé solitons” as solitonic excitations propagating on a background governed by a Painlevé transcendent. A novel symmetry decomposition approach offers a systematic way to derive such solutions [4, 5]. This method leverages the interplay of multiple intrinsic symmetries of the AKNS system. Whereas elliptic solitons arise from the combination of translation invariance and the square eigenfunction symmetry, a different combination-involving scaling invariance and Galilean invariance alongside the square eigenfunction symmetry-leads to a different class of solutions. For the AKNS system encompassing the NLS equation, this specific symmetry combination generates solutions known as “Painlevé IV solitons.”
@@ 2 Symmetry Decomposition Method
$\sigma=\left(\begin{array}[]{l}P-Xp_{x}-Tp_{t}\\ Q-Xq_{x}-Tq_{t}\\ \Phi-X\phi_{x}-T\phi_{t}\\ \Psi-X\psi_{x}-T\psi_{t}\\ F-Xf_{x}-Tf_{t}\end{array}\right),$ (20)
where $P=-ibaxp+m\psi^{2}+\phi_{0}p,\ Q=ibaxq-m\phi^{2}-2cq-\phi_{0}q,\ \Phi=\frac{1}{2}\phi(ibax-2i\sqrt{b}mf-2c-\phi_{0}+n),\ \Psi=-\frac{1}{2}\psi(ibax+2i\sqrt{b}mf-\phi_{0}-n),\ F=-i\sqrt{b}mf^{2}+nf+f_{0},\ X=at+cx+x_{0},\ T=2ct+t_{0}$ with arbitrary constants $a,\ c,\ x_{0},\ t_{0},\ f_{0},\ \lambda,\ m,\ n$ and $\phi_{0}$.
@@ 3 Explicit Solutions and Results
3.1 Elliptic Soliton Solution
@@ 3.2 Painlevé IV Solitons and Irrational Algebraic Solitons
For some special $\alpha$ and $\beta$, solutions of the Painlevé IV equation can be expressed by some types algebraic rational solutions, the special Whittaker functions and/or the parabolic cylinder function $D_{\nu}(\zeta)$ which satisfies
@@ 4 Discussion and Conclusion
We have developed a novel symmetry decomposition approach that reveals how different combinations of intrinsic symmetries of the AKNS system lead to distinct classes of solutions. While the well-established elliptic solitons arise from translation invariance and square eigenfunction symmetry [2], we have discovered that a fundamentally different symmetry combination-scaling invariance, Galilean invariance, and square eigenfunction symmetry-generates Painlevé IV solitons. This represents a major theoretical insight into the structure of integrable systems.
The explicit construction of Painlevé solitons provides exact models for solitary waves interacting with dynamically evolving, non-periodic backgrounds governed by Painlevé transcendents [3, 6]. These solutions significantly expand the known solution landscape of the NLS equation, one of the most important integrable models in mathematical physics. The discovery of irrational algebraic solitons is particularly noteworthy, as these represent a previously unknown class of solutions distinct from conventional rational rogue waves [5].
@@ Acknowledgments
This work was sponsored by the National Natural Science Foundation of China (Grants No. 12235007, 12275144, 12271324, 12301315 and No. 12375003). The authors thank Professors Q. P. Liu and X. B. Hu for helpful discussions.

### v2

@@ 1 Introduction
The study of integrable nonlinear partial differential equations (PDEs) has been a central theme in mathematical physics for over half a century. Among the most influential discoveries in this field is the inverse scattering transform (IST), a powerful method for solving certain nonlinear evolution equations. A major breakthrough in the systematic development of the IST was the formulation of the Ablowitz-Kaup-Newell-Segur (AKNS) system [1, 2]. This elegant framework provides a unified approach to treat a wide class of integrable equations, including the celebrated nonlinear Schrödinger (NLS) [3, 4], Korteweg-de Vries (KdV) [5, 6, 7], modefied KdV [8, 9, 10] and sine-Gordon equations [11, 12, 13], by associating them with a compatible pair of linear differential equations (a Lax pair). Within this scheme, the search for special solutions, particularly solitons, and the investigation of their interactions have revealed profound connections with other branches of mathematics, such as the theory of special functions and Painlevé transcendents.
@@
A significant step in this direction is the concept of the “elliptic solitons” [14]. Here, the soliton propagates on a background given by an elliptic (or cnoidal) wave, which itself is an exact periodic solution of the NLS equation. The elliptic wave is expressed in terms of Jacobi elliptic functions. The composite solution, the elliptic solitons, can be derived through sophisticated methods such as the Darboux transformation or Bäcklund transformation, applied to the finite-gap (periodic or quasi-periodic) solutions of the integrable system. Ling and Sun [15] construct multiple elliptic solitons and analyze their asymptotic behaviors for the focusing modified KdV equation via Darboux-Bäcklund transformation. Nijhoff, Sun and Zhang [16] introduced the concept of elliptic $N$th root of unity and established an elliptic direct linearization scheme (of Dutch version). Elliptic $N$th root of unity can be used to defined plane wave factors for the Boussinesq equations and the equations in the Gal’fand-Dicky hierarchy. A bilinear framework serving for elliptic solitons was established by Li and Zhang for both continuous [17] and discrete systems. More recently, an elliptic direct linearization scheme (Fokas-Ablowitz’s version) was established by Li, Sun and Zhang [18] for the KP, KdV and Boussinesq equations. The existence of such solutions is deeply tied to the underlying symmetries of the equation. Specifically, it has been shown that the combination of the space-time translation invariance and the so-called nonlocal symmetries [19] such as the pseudopotential symmetries [20], residual symmetries [21], square eigenfunction symmetry [22, 23] is instrumental in generating these elliptic soliton solutions.
@@
where $\alpha$ and $\beta$ are complex parameters, emerges naturally in similarity reductions of the NLS equations, the derivative NLS equations, the Boussinesq equation [25, 26] and other integrable systems. Its solutions, the Painlevé IV transcendent, provide a rich class of potential backgrounds that are asymptotically rational, algebraic, or expressible in terms of parabolic cylinder functions and error functions.
By analogy with elliptic solitons, one can define “Painlevé solitons” as solitonic excitations propagating on a background governed by a Painlevé transcendent. A novel symmetry decomposition approach offers a systematic way to derive such solutions [26]. This method leverages the interplay of multiple intrinsic symmetries of the AKNS system. Whereas elliptic solitons arise from the combination of translation invariance and the square eigenfunction symmetry, a different combination-involving scaling invariance and Galilean invariance alongside the square eigenfunction symmetry-leads to a different class of solutions[26]. For the AKNS system encompassing the NLS equation, this specific symmetry combination generates solutions known as “Painlevé IV solitons.”
@@ 2 Symmetry Decomposition Method
$\sigma\equiv\left(\begin{array}[]{l}\sigma^{p}\\ \sigma^{q}\\ \sigma^{\phi}\\ \sigma^{\psi}\\ \sigma^{f}\\ \sigma^{\lambda}\end{array}\right)=\left(\begin{array}[]{l}P-Xp_{x}-Tp_{t}\\ Q-Xq_{x}-Tq_{t}\\ \Phi-X\phi_{x}-T\phi_{t}\\ \Psi-X\psi_{x}-T\psi_{t}\\ F-Xf_{x}-Tf_{t}\\ \delta\end{array}\right),$ (20)
where
$P$ $=-ibaxp+m\psi^{2}+\phi_{0}p,$ (21)
$Q$ $=ibaxq-m\phi^{2}-2cq-\phi_{0}q,$
$\Phi$ $=\frac{1}{2}\phi(ibax-2i\sqrt{b}mf-2c-\phi_{0}+n),$
$\Psi$ $=-\frac{1}{2}\psi(ibax+2i\sqrt{b}mf-\phi_{0}-n),$
$F$ $=-i\sqrt{b}mf^{2}+nf+f_{0},$
$X$ $=at+cx+x_{0},\ T=2ct+t_{0},\ \delta=-c\lambda-\frac{1}{2}ab$
with arbitrary constants $a,\ c,\ x_{0},\ t_{0},\ f_{0},\ \lambda,\ m,\ n$ and $\phi_{0}$.
A Lie point symmetry (20) with (21) of the extended AKNS system implies that (10)–(17) and (3) are invariant under the transformation
$\{p,\ q,\ \phi,\ \psi,\ f,\ \lambda\}\rightarrow\{p,\ q,\ \phi,\ \psi,\ f,\ \lambda\}+\epsilon\{\sigma^{p},\ \sigma^{q},\ \sigma^{\phi},\ \sigma^{\psi},\ \sigma^{f},\ \delta\},$
with infinitesimal parameter $\epsilon$.
@@ 3 Explicit Solutions and Results
3.1 Elliptic Soliton Solutions
@@ 3.2 Painlevé IV Solitons and Irrational Algebraic Solitons
For some special $\alpha$ and $\beta$, solutions of the Painlevé IV equation can be expressed by some types algebraic rational solutions, the special Whittaker functions and/or the parabolic cylinder function ${\cal{D}}_{\nu}(\zeta)$ which satisfies
@@
Because the $F$-equation (28) and the Painlevé IV equation (2) are related via (32), the argument of the Painlevé transcendent becomes complex with respect to the real space-time variables. Consequently, if we start directly from the solutions of the Painlevé IV equation, we can readily obtain solutions of the AKNS system, but not those of the NLS equation, since the complex conjugate condition $q=p^{*}$ is difficult to satisfy.
It is not difficult to verified that the Painlevé IV equation (2) possesses a special solution
$w=\frac{(2\nu+1){\cal{D}}_{\nu+1}(\sqrt{2}z)}{\sqrt{2}{\cal{D}}_{\nu}(\sqrt{2}z)},\ \alpha=\left(\nu-\frac{1}{2}\right)\delta,\ \beta=-\frac{1}{2}(2\nu+1)^{2},\ \delta^{2}=1,$ (39)
where ${\cal{D}}_{\nu}(\zeta)$ is the parabolic cylinder function which is a solution of (33).
Correspondingly, the $F$-function takes the form (with $\delta^{2}=\delta_{1}^{2}=\delta_{2}^{2}=\delta_{3}^{2}=1$)
$F=\frac{4\delta}{2\nu+1}\ln\left[\mathcal{D}_{\nu+1}\left(\frac{\sqrt{2}\,\delta(\delta_{1}+\mathrm{i}\delta_{2})}{2}\,\eta\right)\right],\quad A=\frac{\delta_{3}}{8}(2\nu+1),\quad\phi_{0}=\frac{1}{2}b\delta\delta_{1}\delta_{2}(2\nu-1)-1,$ (40)
and the corresponding solution of the AKNS system is given by (27) together with (40). Figure 4 displays the parabolic cylinder soliton of the AKNS system for the quantity $I\equiv pq$, where $p$ and $q$ are given by (27) and $F$ is specified as in (40) with the parameter choices $\delta=\delta_{1}=\delta_{2}=\delta_{3}=-b=1$, $\lambda=0$, and $\nu=2$.
As shown in Fig. 4(b), the imaginary part of $pq$ is nonzero. Therefore, the corresponding solution belongs only to the AKNS system and does not satisfy the NLS equation.
Figure 4: Structure of the parabolic cylinder soliton of the AKNS system for the quantity $I\equiv pq$, where $p$ and $q$ are given by (27) and $F$ is specified as in (40) with the parameter choices $\delta=\delta_{1}=\delta_{2}=\delta_{3}=-b=1$, $\lambda=0$, and $\nu=2$: (a) Real part and (b) imaginary part of II.
@@ 4 Discussion and Conclusion
We have developed a novel symmetry decomposition approach that reveals how different combinations of intrinsic symmetries of the AKNS system lead to distinct classes of solutions. While the well-established elliptic solitons [14, 15, 16] arise from translation invariance and square eigenfunction symmetry [20, 31, 32], we have discovered that a fundamentally different symmetry combination-scaling invariance, Galilean invariance, and square eigenfunction symmetry-generates Painlevé IV solitons. This represents a major theoretical insight into the structure of integrable systems.
The explicit construction of Painlevé solitons provides exact models for solitary waves interacting with dynamically evolving, non-periodic backgrounds governed by Painlevé transcendents [24, 27]. These solutions significantly expand the known solution landscape of the NLS equation, one of the most important integrable models in mathematical physics. The discovery of irrational algebraic solitons is particularly noteworthy, as these represent a previously unknown class of solutions distinct from conventional rational rogue waves [33, 34].
@@ Acknowledgments
This work was sponsored by the National Natural Science Foundation of China (Grants No. 12235007, 12275144, 12271324, 12301315 and No. 12375003). The authors thank Professors Q. P. Liu, X. B. Hu, B. F. Feng, Y. R. Xia and Dr. Y. Li for their helpful discussions.